# Learning Molecular Representation in a Cell

### v2

@@ Abstract
Predicting drug efficacy and safety in vivo requires information on biological responses (e.g., cell morphology and gene expression) to small molecule perturbations. However, current molecular representation learning methods do not provide a comprehensive view of cell states under these perturbations and struggle to remove noise, hindering model generalization. We introduce the Information Alignment (InfoAlign) approach to learn molecular representations through the information bottleneck method in cells. We integrate molecules and cellular response data as nodes into a context graph, connecting them with weighted edges based on chemical, biological, and computational criteria. For each molecule in a training batch, InfoAlign optimizes the encoder’s latent representation with a minimality objective to discard redundant structural information. A sufficiency objective decodes the representation to align with different feature spaces from the molecule’s neighborhood in the context graph. We demonstrate that the proposed sufficiency objective for alignment is tighter than existing encoder-based contrastive methods. Empirically, we validate representations from InfoAlign in two downstream tasks: molecular property prediction against up to 19 baseline methods across four datasets, plus zero-shot molecule-morphology matching.
@@ 1 Introduction
Drug properties, e.g., toxicity and adverse effects [25], are induced by molecular initiating events—interactions between a molecule and a biological system—that first impact the cellular level and ultimately influence tissue or organ functions [31, 25]. However, a chemical molecule’s structure alone is insufficient information to predict its impact on cells: each chemical interacts with multiple cells and genes and induces complex changes in gene expression and cell morphology, making predictions of downstream responses challenging [5, 49, 32]. Hence, molecular representation learning should make use of information about cellular response, enhancing the representation of the mode of action and thereby improving predictions for downstream bioactivity tasks [25, 53].
@@
Figure 1: Molecular Representation Learning via the Information Bottleneck: (a) Existing contrastive learning methods utilize two encoders—one for molecules and another for cell morphology or gene expression features, lacking a holistic view of molecular representation learning in cells. (b) In contrast, InfoAlign excludes redundant information in molecule-cell morphology or molecule-gene expression contrastive pairs that could impair generalization [53] for bottleneck representation.
Cells can be perturbed by treating them with chemicals or genetic reagents that disrupt a particular gene or pathway. These chemical and genetic perturbations in vitro naturally bridge molecules with cell morphology and gene expression, as illustrated in Figure 1 (b). However, multi-modal contrastive methods such as CLOOME [45] and InfoCORE [53], depicted in Figure 1 (a), focus primarily on aligning molecular representations with cell morphology [45, 53] or gene expression [53]. These approaches fall short in two ways. (1) They do not remove redundant information, grey-colored area in Figure 1 (b), that may harm representation generalization. The presence of redundant information [53] may induce spurious correlations, adversely affecting the generalization of molecular representations. For example, in small molecule perturbations [3, 6], batch identifiers can signify confounding technical factors, creating misleading associations between molecular structures and cell morphology [53]. (2) They treat molecules as the sole connectors between gene expression and cell morphology, ignoring the potential for genetic perturbations [6] to directly establish connections between these modalities. The data derived from genetic perturbations [6] not only enriches the feature space of gene expression and cell morphology but also enhances the navigation of molecular representation learning towards the overlapped (bottleneck) area in Figure 1 (b).
To address the aforementioned challenges, we conceptualize the cellular response processes as a context graph, capturing a more complete set of interactions among molecules, gene expression, and cell morphology. Using the context graph, we identify neighborhoods for each molecule and apply the information bottleneck [52] to optimize molecular representations, which aligns them with neighboring biological variables to remove redundant information and improve generalization.
@@
The sufficiency objective introduces a decoder-based bound for multi-modal alignment using decoders. We demonstrate the theoretical advantages of this approach, showing it provides a tighter bound than the encoder-based bound used in previous contrastive methods [35, 39, 53, 45] in Section 4.3. In experiments, InfoAlign outperforms up to 19 representation learning baselines across three classification and one regression datasets, with overall improvements of +10.58% and +6.33% on the Broad6K classification and Biogen3K regression datasets, respectively. InfoAlign also exhibits highly competitive zero-shot multi-modal matching capabilities on two molecule-morphology datasets.
@@ 2 Related Work
Representation learning approaches for molecules can be categorized into sequential-based [24, 44] or graph-based models [17, 57, 58, 27]. Sequential models, utilizing string formats of molecules like SMILES and SELFIES [24], have evolved from Recurrent Neural Networks (RNNs) to Transformers [7, 44]. These models typically follow specific pretraining strategies similar to language models such as BERT [9], RoBERTa [30, 7] and GPT [38]. The pretraining targets are thus often the next token predictions or mask language modeling [9, 7] on SMILES or SELFIES sequences [38]. Graph Neural Networks (GNNs) are the architectures for graph-based approaches [17, 57, 58, 29], where methods to pretrain GNNs often perturb or mask the atoms, edges, or substructures of molecular graphs with contrastive [17, 57] and predictive losses [58, 20]. Recent evidence highlights the challenges of developing universal molecular representations based solely on molecular structures without integrating domain knowledge [3, 46, 50, 47, 28]. Although using motifs is a common method to incorporate such knowledge [43, 20], the incorporation of information about molecules’ biological impacts is much less explored. We aim to enhance molecular representation learning by incorporating domain knowledge from cellular response data.
Representation Learning with Different Modalities:
@@
Existing methods on multimodal alignment, such as CLIP [39], primarily address pairwise relationships between texts and images and use methods like InfoNCE [35, 53, 45]. These approaches use separate encoders for different modalities to compute contrastive loss, which is upper bounded by the number of negative examples [37]. Subsequent research on molecules similarly focuses on pairwise alignment between molecules and cell images [45, 53], molecules and protein sequences [19], and molecules and text [10, 21]. Although BioBridge [55] handles multiple modalities, it leverages a knowledge graph for transforming representations between modalities rather than optimizing molecular representations.
Representation Learning with Cellular Response Data:
@@
A primary goal of molecular representation learning is to predict molecular bioactivity. Likewise, emerging gene expression [49] and morphological profiling approaches [5, 3, 6, 48] that describe perturbed genetic or cellular states in cell cultures can also be used to predict bioactivity. In some datasets, molecules are the perturbations, and the perturbed cell states measured are gene expression values for a thousand or more genes [49] and/or microscopy Cell Painting images, which can be represented as a thousand or more morphology features [8]. Recently created large-scale perturbation datasets [3, 49, 6] could enrich molecular representation learning approaches. CLOOME [45] and MoCoP [33] contrast cellular images with molecules and InfoCORE [53] contrasts molecule with either morphological profiling [4] or gene expression [53]. InfoCORE [53] aims to mitigate confounding batch identifiers, but its effectiveness depends on a batch classifier, which is impractical without batch identifiers. We integrate cellular response data and molecules into a context graph to capture cellular response patterns, focusing on learning molecular representations to remove nuisances [51].
@@ 3 Problem Definition
Suppose a molecule $X\in\mathcal{X}$ is from the space $\mathcal{X}$ constructed by atoms and bonds. An encoder with parameter $\theta$ could map $X$ into a $D$-dimensional representation space $\mathcal{Z}$: $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Z}\subset\mathbb{R}^{D}$. The encoder $f_{\theta}$ could be learned with Graph Neural Networks (GNNs) [23, 56], where atoms are nodes and bonds are edges. Such GNNs first update node representations and then perform a readout operation (e.g., summing) over nodes to obtain the latent representation variable $Z\in\mathcal{Z}$ for $X$.
Existing research in contrastive or generative learning has extensively leveraged structural features to pretrain the GNN encoder $f_{\theta}$ [17, 20, 28]. However, using more expressive features from the cellular context, such as cell morphology $C$ and gene expression $E$, remains largely unexplored for improving molecular representations.
4 The Proposed InfoAlign
Figure 2: Representation Learning Over Walk Paths in Context Graphs: (a) In Section 4.1, we construct the graph with various interaction, perturbation, and cosine similarities among molecules $X$, cell morphology profiles $C$, and gene expression profiles $E$. Given a training batch of molecules, including $X_{1}$ and $X_{4}$, random walk extracts paths, for instance, of length 4. (b) In Section 4.2, we aim to learn bottleneck representations from molecules with encoder $f_{\theta}$ that minimally relate to structures yet maximally reconstruct feature with decoders $g_{\phi}$ about variables along the walk paths.
@@ 4.1 Random Walks on Cellular Context Graph
In the context graph, molecules, cell morphology, and genes with expression values serve as nodes with feature spaces $\mathcal{Y}^{[X]}$, $\mathcal{Y}^{[C]}$, and $\mathcal{Y}^{[E]}$ for vectors and scalars, respectively. Molecular features are vectors obtained using fingerprint [42]. Cell morphology features are vectors derived from CellProfiler [5] applied to Cell Painting microscopy images. Gene expression features are scalars using L1000 [49] methods. We further rescale the feature spaces to a range between 0 and 1.
We link nodes using various chemical, biological, and computational criteria. For example, molecules can perturb cultured human cells, inducing changes in cell morphology [6] and gene expression [49], thus linking them to cell morphology and gene expression nodes. Genes could also perturb cells, inducing links between genes and cell morphology [6]. Additionally, we calculate cosine similarity within the same feature space $\mathcal{Y}$ and use biological criteria such as gene-gene interactions [16] to enrich the edge space. Formally, we define:
Definition 4.1 (Cellular Context Graph).
The graph $G=(\mathcal{V},\mathcal{E})$ comprises nodes $v\in\mathcal{V}$, each denoting a data point and mapped to a feature space $\psi:\mathcal{V}\rightarrow\{\mathcal{Y}^{[X]},\mathcal{Y}^{[C]},\mathcal{Y}^{[E]}\}$. Each edge $e\in\mathcal{E}$ has a weight from 0 to 1, indicating the similarity between nodes based on specific criteria.
We provide more details on the data sources, node features, and edge weights for constructing the context graph in Section 5.1.
Path Extraction with Random Walk
@@
4.2 Optimization for Bottleneck Representation
@@
We aim to extract minimal sufficient information from $X$ through aligning its representation with features from variables along the walk path $\mathcal{P}_{X}$, which could effectively mitigate nuisance [51]. To achieve this, we define the bottleneck representation of molecule $X$ using two mutual information (MI) principles: (1) the minimality principle, which minimizes MI between molecules and their latent representations as $I(X;Z)$, and (2) the sufficiency principle, which decodes latent representations to maximally reconstruct feature spaces for variables along the walk path $\sum_{v\in\mathcal{P}_{X}}I\left(Z;\psi(v)\right)$. Together, these form the minimal sufficient optimization objectives:
$\min_{p(Z|X)}\left[-\sum_{v\in\mathcal{P}_{X}}I\left(Z;\psi(v)\right)+\beta I(X;Z)\right],$ (1)
@@
where $\beta$ controls the tradeoff between minimality and sufficiency. The exact computation of $I(Z;\psi(v))$ and $I(X;Z)$ is intractable due to the unknown conditional distribution $p(\psi(v)|Z)$ and the marginal $p(Z)$. We introduce the variational approximations $q(\psi(v)|Z)$ and $q(Z)$ for them, respectively. This results in a lower bound estimation for the first decoding term $I_{DLB}$ and an upper bound for the second encoding term $I_{EUB}$ [37].
$I(Z;\psi(v))$ $\geq\mathbb{E}_{p(Z,\psi(v))}\left[\log q(\psi(v)|Z)\right]+H(\psi(v))\triangleq I_{DLB}$ (2)
$I(X;Z)$ $\leq\mathbb{E}_{p(X)}\left[\operatorname{KL}\left(p(Z|X)\|q(Z)\right)\right]\triangleq I_{EUB}$
$H(\psi(v))$ is the differential entropy. Proofs are in Section A.1. Combined together, $I_{DLB}$ and $I_{EUB}$ upper bound the Eq. 1 to form a tractable objective $-I_{DLB}+I_{EUB}$ to optimize the encoder $f_{\theta}$ with different feature decoders $g_{\phi}^{[X]},g_{\phi}^{[E]},g_{\phi}^{[C]}$ for molecular features, gene expression features, and cell morphology features, respectively. Specifically, the encoder $f_{\theta}(X)$ learns the mean and variance $\boldsymbol{\mu}_{Z}\in\mathbb{R}^{D}$, $\boldsymbol{\sigma}_{Z}\in\mathbb{R}^{D}$ of the latent variable $Z$. The decoder $g_{\phi}$ could take as inputs the mean to reconstruct features for the neighboring variables of $X$ on the context graph. After ignoring the constant terms, we could implement Eq. 2 as the loss function for the molecule $X$:
@@
$\mathcal{L}=\frac{1}{L}\sum_{v\in\mathcal{P}_{X}}\alpha(v)\operatorname{CE}\left(g_{\phi}^{[v]}(f_{\theta}(X)),\psi(v)\right)+\beta\operatorname{KL}\left(f_{\theta}(X)\|\mathcal{N}(0,I)\right),$ (3)
where $\operatorname{CE}$ and $\operatorname{KL}$ are cross entropy and Kullback–Leibler divergence, respectively. The Gaussian $\mathcal{N}(0,I)$ is the prior distribution [1]. With reparameterization [22], we can directly backpropagate through the neural network to optimize parameters $\theta$ and $\phi$.
@@
InfoAlign uses multiple decoders $g_{\phi}$ to align multi-modal features, while prior work relies on multi-encoders with CLIP-like losses to align the latent space [39, 14, 53, 45]. Next, we provide the theoretical benefits of decoder-based alignment alongside the empirical advantages in Section 5.

### v3

@@ Abstract
Predicting drug efficacy and safety in vivo requires information on biological responses (e.g., cell morphology and gene expression) to small molecule perturbations. However, current molecular representation learning methods do not provide a comprehensive view of cell states under these perturbations and struggle to remove noise, hindering model generalization. We introduce the Information Alignment (InfoAlign) approach to learn molecular representations through the information bottleneck method in cells. We integrate molecules and cellular response data as nodes into a context graph, connecting them with weighted edges based on chemical, biological, and computational criteria. For each molecule in a training batch, InfoAlign optimizes the encoder’s latent representation with a minimality objective to discard redundant structural information. A sufficiency objective decodes the representation to align with different feature spaces from the molecule’s neighborhood in the context graph. We demonstrate that the proposed sufficiency objective for alignment is tighter than existing encoder-based contrastive methods. Empirically, we validate representations from InfoAlign in two downstream applications: molecular property prediction against up to 27 baseline methods across four datasets, plus zero-shot molecule-morphology matching.
@@ 1 Introduction
Drug properties, e.g., toxicity and adverse effects [25], are induced by molecular initiating events—interactions between a molecule and a biological system—that first impact the cellular level and ultimately influence tissue or organ functions [32]. However, a chemical molecule’s structure alone is insufficient information to predict its impact on cells: each chemical interacts with multiple cells and genes and induces complex changes in gene expression and cell morphology, making predictions of downstream responses challenging [5, 33]. Hence, molecular representation learning should make use of information about cellular response, enhancing the representation of the mode of action and thereby improving predictions for downstream bioactivity tasks [25, 54].
@@
Figure 1: Comparison of Representation Learning Methods: (a) Existing contrastive methods use two encoders—one for molecules and another for cell morphology or gene expression features—without sharing the molecule encoders for different alignment targets. (b) InfoAlign remove redundant information from molecules, cell morphology, and gene expressions based on the information bottleneck, resulting in more concise yet predictive molecular representations [1].
Cells can be perturbed by treating them with chemicals or genetic reagents that disrupt a particular gene or pathway. These chemical and genetic perturbations in vitro naturally bridge molecules with cell morphology and gene expression, as illustrated in Figure 1 (b). However, multi-modal contrastive methods such as CLOOME [46] and InfoCORE [54], depicted in Figure 1 (a), focus primarily on aligning molecular representations with cell morphology [46, 54] or gene expression [54]. These approaches fall short in two ways. (1) They do not remove redundant information, grey-colored area in Figure 1 (b), that may harm representation generalization. The presence of redundant information [54] may induce spurious correlations, adversely affecting the generalization of molecular representations. For example, in small molecule perturbations [3, 6], batch identifiers can signify confounding technical factors, creating misleading associations between molecular structures and cell morphology [54]. (2) They treat molecules as the sole connectors between gene expression and cell morphology, ignoring the potential for genetic perturbations [6] to directly establish connections between these modalities. Genetic perturbations [6] not only enrich the feature space of gene expression and cell morphology but also enhance the navigation of molecular representation learning towards the overlapped (bottleneck) area in Figure 1 (b).
To address the aforementioned challenges, we conceptualize the cellular response processes as a context graph, capturing a more complete set of interactions among molecules, gene expression, and cell morphology. We identify the neighborhood of the molecule on the context graph and apply the information bottleneck [53] to optimize molecular representations, which aligns them with neighboring biological variables to remove redundant information and improve generalization.
@@
The sufficiency objective introduces a decoder-based bound for multi-modal alignment. We show its theoretical advantages by demonstrating that it provides a tighter bound than the encoder-based approaches used in previous contrastive methods [36, 40], as discussed in Section 4.3. In experiments, InfoAlign outperforms up to 27 baselines across three classification and one regression dataset, covering 685 tasks, with average improvements of up to 6.4%. InfoAlign also demonstrates strong zero-shot multi-modal matching on two molecule-morphology datasets.
@@ 2 Related Work
Representation Learning on Molecular Structure: Representation learning approaches for molecules can be categorized into sequential-based [24, 45] or graph-based models [18, 60, 62, 27]. Sequential models, utilizing string formats of molecules like SMILES and SELFIES [24], have evolved from Recurrent Neural Networks (RNNs) to Transformers [7, 45]. These models typically follow specific pretraining strategies similar to language models such as BERT [9], RoBERTa [30, 7] and GPT [39]. The pretraining targets are thus often the next token predictions or mask language modeling [9, 7] on SMILES or SELFIES sequences [39]. Graph Neural Networks (GNNs) are the architectures for graph-based approaches [18, 60, 62, 29], where methods to pretrain GNNs often perturb or mask the atoms, edges, or substructures of molecular graphs with contrastive [18, 60] and predictive losses [62, 21]. Recent evidence highlights the challenges of developing universal molecular representations based solely on molecular structures without integrating domain knowledge [3, 47, 51, 48, 28]. Although using motifs is a common method to incorporate such knowledge [44, 21], the incorporation of information about molecules’ biological impacts is much less explored. We aim to enhance molecular representation learning by incorporating domain knowledge from cellular response data.
@@
Representation Learning with Different Modalities: Existing methods on multimodal alignment, such as CLIP [40], primarily address pairwise relationships between texts and images and use methods like InfoNCE [36, 54, 46]. These approaches use separate encoders for different modalities to compute contrastive loss, which is upper bounded by the number of negative examples [38]. Subsequent research on molecules similarly focuses on pairwise alignment between molecules and cell images [46, 54], molecules and protein sequences [20], and molecules and text [10, 22]. Although BioBridge [57] handles multiple modalities, it leverages a knowledge graph for transforming representations between modalities rather than optimizing molecular representations.
@@
Representation Learning with Cellular Response Data: A primary goal of molecular representation learning is to predict molecular bioactivity. Likewise, emerging gene expression [50] and morphological profiling approaches [5, 49] that describe perturbed genetic or cellular states in cell cultures can also be used to predict bioactivity. In some datasets, molecules are the perturbations, and the perturbed cell states measured are gene expression values for a thousand or more genes [50] and/or microscopy Cell Painting images, which can be represented as a thousand or more morphology features [8]. Recently created large-scale perturbation datasets [3, 50, 6] could enrich molecular representation learning approaches. CLOOME [46] and MoCoP [34] contrast cellular images with molecules and InfoCORE [54] contrasts molecule with either morphological profiling [4] or gene expression [54]. InfoCORE [54] aims to mitigate confounding batch identifiers, but its effectiveness depends on a batch classifier, which is impractical without batch identifiers. We integrate cellular response data and molecules into a context graph to capture cellular response patterns, focusing on learning molecular representations to remove nuisances [52].
@@ 3 Problem Definition
We denote $x\in\mathcal{X}$ as the molecule from the space $\mathcal{X}$. An encoder with parameters $p_{\theta}(\mathbf{z}\mid x)$ maps $x$ to a $D$-dimensional latent representation $\mathbf{z}\in\mathbb{R}^{D}$. One may implement a Graph Neural Network (GNN) [58] as the encoder. The GNN first updates node representations and then performs a readout operation (e.g., summation) over the nodes to obtain the latent representation.
Existing research has extensively used structural features to pretrain the GNN encoder [18, 21]. However, incorporating more expressive features from the cellular context, such as cell morphology and gene expression, remains largely unexplored for improving molecular representations. In this work, we use these features as targets to optimize molecular representations.
4 Multi-modal Alignment with InfoAlign
Figure 2: Molecular Representation Learning Using the Context Graph: (a) In Section 4.1, we construct the graph with various interaction, perturbation, and cosine similarities among molecules $x$, cell morphology profiles $c$, and genes $e$. Given a training batch of molecules, including $x_{1}$ and $x_{4}$, random walk extracts paths, for instance, of length four. (b) In Section 4.2, we aim to learn molecular representations based on the information bottleneck, preserving minimal information from the input molecule while ensuring sufficient information for decoding the target along the walk path $\mathcal{P}_{x}$.
@@ 4.1 Random Walks on Cellular Context Graph
Node Construction: We model the interactions of the molecule $x$ with other molecules, the cell $c$, and genes $e$ using the context graph. They are nodes with different features $y$. Molecular features are vectors obtained using fingerprint [43]. Cell morphology features are vectors derived from CellProfiler [5] applied to Cell Painting microscopy images. Gene expression features are scalars using L1000 [50] methods. We further rescale the feature spaces to a range between 0 and 1.
Edge Construction: We link nodes using various chemical, biological, and computational criteria. For example, molecules can perturb cultured human cells, inducing changes in cell morphology [6] and gene expression [50], thus linking them to cell morphology and gene expression nodes. Genes could also perturb cells, inducing links between genes and cell morphology [6]. Additionally, we calculate cosine similarity within the same feature space and use biological criteria such as gene-gene interactions [16] to enrich the edge space. Each edge is assigned a weight $w$ with a value between 0 and 1. We construct the context graph with details Section 5. An example is provided in Figure 2 (a).
@@
4.2 Optimization for Representation with Information Bottleneck
@@
The information bottleneck (IB) [53, 1] is an appealing method for defining concise representations with strong predictive power. For molecular representation, we extract minimal sufficient information from the random variable $X$ of molecules. This is achieved by aligning the molecular representations $Z$ with the targets $Y$, derived from node features along the walk path $\mathcal{P}$. The IB has two principles based on mutual information (MI): (1) the minimality principle, which minimizes MI between molecules and their latent representations as $I(X;Z)$, and (2) the sufficiency principle, which decodes latent representations to maximally reconstruct feature spaces for variables along the walk path $I\left(Z;Y\right)$. Together, these form the optimization objectives:
$\min_{p(\mathbf{z}\mid x)}\left[-I\left(Z;Y\right)+\beta I(X;Z)\right],$ (1)
@@
where $\beta$ controls the trade-off between minimality and sufficiency. The exact computation of $I(Z;Y)$ and $I(X;Z)$ is intractable due to the unknown conditional distribution $p(y|\mathbf{z})$ and the marginal $p(\mathbf{z})$. We introduce the variational approximations $q(y|\mathbf{z})$ and $q(\mathbf{z})$ for them, respectively. This results in a lower bound estimation for the first decoding term $I_{DLB}$ and an upper bound for the second encoding term $I_{EUB}$ [38].
$I(Z;Y)$ $\geq\mathbb{E}_{p(\mathbf{z},y)}\left[\log q(y\mid\mathbf{z})\right]+H(Y)\triangleq I_{DLB}$ (2)
$I(X;Z)$ $\leq\mathbb{E}_{p(x)}\left[\operatorname{KL}\left(p(\mathbf{z}\mid x)\|q(\mathbf{z})\right)\right]\triangleq I_{EUB}$
$H(Y)$ is the differential entropy. Proofs are in Section A.1. Together, $I_{DLB}$ and $I_{EUB}$ upper bound Eq. 1, forming a tractable objective $-I_{DLB}+I_{EUB}$ to optimize the encoder. For the target $Y$, the $I_{DLB}$ objective requires decoders rather than encoders, as typically used in prior work [46]. We use distinct decoders, denoted as $q_{\phi}$ with parameters $\phi$, for various targets, including molecular fingerprints, gene expressions, and cell morphology features.
After ignoring the constant terms, one could formulate the loss function according to Eq. 2 for the molecule sample $x$, its latent representation $\mathbf{z}$, and the targets $y_{v}$ from $\mathcal{P}_{x}$:
@@
$\mathcal{L}=\frac{1}{L}\sum_{v\in\mathcal{P}_{x}}\alpha(v|\mathcal{P}_{x})\left[-\log\left(q_{\phi}(y_{v}\mid\mathbf{z}\right)\right]+\beta\operatorname{KL}\left(p_{\theta}(\mathbf{z}\mid x)\mid\mathcal{N}(0,I)\right),$ (3)
where the first term aligns the representation with other features, and $\operatorname{KL}$ is the Kullback–Leibler divergence used for regularization. $\mathcal{N}(0,I)$ is the The Gaussian prior. In this formulation, the encoder models a distribution instead of a single representation $\mathbf{z}$, learning the mean and variance $\boldsymbol{\mu},\boldsymbol{\sigma}\in\mathbb{R}^{D}$. One may use parameterization tricks to sample $\mathbf{z}$ from the distribution [1]. The decoder then reconstructs $y_{v}$, the features of the neighboring node $v$ on the context graph.
@@
InfoAlign uses multiple decoders for $q_{\phi}$ to align multi-modal features, while prior work relies on encoders with CLIP-like losses to align the latent space [40, 14, 54, 46]. Next, we provide the theoretical benefits of decoder-based alignment alongside the empirical advantages in Section 6.